Suppose that earning a college degree will increase your annual salary by $24,000.00. If you have just finished this degree and plan on working for the next 44.00 years before retiring, what is the present value of what the degree will earn you across your entire career? Assume a market interest rate of 8.00%. Round your answer to two decimals.



1
Expert's answer
2021-10-22T08:16:16-0400

"PV= PMT\times\frac{1-(\frac{1}{(1+r)^n})}{r}"


"PV=24,000\times\frac{1-(\frac{1}{(1+0.08)^{44}})}{0.08}"


PV= $289,849.77